### v3

# Multimodal Visual Transformer for Sim2real Transfer in Visual Reinforcement Learning

Abstract
Depth information is robust to scene appearance variations and inherently carries 3D spatial details. In this paper, a visual backbone based on the vision transformer is proposed to fuse RGB and depth modalities for enhancing generalization. Different modalities are first processed by separate CNN stems, and the combined convolutional features are delivered to the scalable vision transformer to obtain visual representations. Moreover, a contrastive unsupervised learning scheme is designed with masked and unmasked tokens to accelerate the sample efficiency during the reinforcement learning process. Simulation results demonstrate that our visual backbone can focus more on task-related regions and exhibit better generalization in unseen scenarios. For sim2real transfer, a flexible curriculum learning schedule is developed to deploy domain randomization over training processes. Finally, the feasibility of our model is validated to perform real-world manipulation tasks via zero-shot transfer.
Index Terms:
Visual Reinforcement Learning, Representation Learning, Sim2Real
I INTRODUCTION
Reinforcement learning (RL) has exhibited its superior ability in addressing contact-rich tasks without a tedious dynamics model. Recent works focus on integrating different modality information tailored to specific task scenarios, in which vision [1], proprioception [2], force/torque [3], and tactile [4] are prevalent alternatives. Considering real-world situations and the cost of acquiring information, vision is gradually emerging as an essential modality. In contrast to state-based data, high-dimensional vision streams need to be encoded as compact representations via visual backbones. For this reason, pixel-based RL exhibits low sample efficiency, i.e., more environment steps are required to achieve the same performance level. Thus, several works have been devoted to improving the sample efficiency of visual RL [5, 6].
Nevertheless, the substantial training cost of visual RL still discourages training on real robots. Researchers try to save the training cost via sim2real or fine-tuning in reality [7, 8]. For the exploitation of visual information, a visual backbone pretrained on in-the-wild datasets has been demonstrated to be effective in extracting generalized representations across different conditions, which also contributes to improving sim2real success rates [9, 10]. However, generalizability is a relatively ambiguous metric for evaluating the visual backbone. Some works suggest that a vision backbone capable of extracting object- or agent-centric representations performs better in diverse environments, benefiting from its weak sensitivity to irrelevant distractors [11, 12]. RGB images and domain randomization are employed to narrow the gap between simulation and reality in previous works [13, 14], which is limited to boost the transfer performance. Thus, other modalities are pursued for additional gains.
Compared to RGB images, depth information can better mitigate perceptual disparities and provide distance information to empower spatial awareness. Depth images enable more efficient online training than other forms of depth representations, e.g., point clouds, with less training burden. A common utilization of depth images is early fusion, in which depth is an extra channel of RGB images. Empirical evidence from computer vision demonstrates that the fusion strategy and framework for different modalities play a pivotal role in semantic segmentation or detection [15, 16], revealing that the power of depth information can be further unleashed through reasonable vision encoders. Inspired by the above, this paper introduces a deeper fusion paradigm with the self-attention mechanism of the vision transformer (ViT) [17]. We also compare and analyze multiple approaches for fusing depth information. Extensive simulation experiments demonstrate that the proposed vision backbone can better utilize the depth information and thus improve the sample efficiency and generalization. In summary, our contributions are threefold:
Figure 1: Zero-shot transfer to complete manipulation tasks.
  1. 1.
    We propose a lightweight backbone based on ViT to fuse RGB-D information. The trained vision backbone demonstrates the capability to extract object- and agent-centric representations while effectively exploiting the complementary advantages of each modality.
  2. 2.
    The contrastive learning scheme is configured with masked and unmasked convoluted features, which are encoded by ViT to obtain query and key latents.
  3. 3.
    The visual backbone and policy network trained through curriculum-based domain randomization can be transferred to the real world in a zero-shot manner.
II RELATED WORK
II-A Visual RL
Visual RL is an intersection field spanning RL and computer vision. A prevalent framework is to integrate the training of the visual backbone with the policy network, which facilitates the extraction of in-domain features [18, 19]. For this training paradigm, limited visual diversity and unlabeled data highlight the importance of improving sample efficiency. Data augmentations [20], such as random shift and color jitter, can significantly enhance the sample efficiency and adaptability of the agent to unseen environments. Extra recipes are essential for stabilizing critic updates during training. Denis et al. proposed the renowned DrQ [21] and its successor DrQv2 [22], which derive the critic target from weak augmented visual observations. While SVEA [23] allows weak and strong augmented observations to share the same critic target. In addition to meticulous recipes, visual backbone is also relevant to sample efficiency and visual attention regions. A CNN-based visual backbone is more preferred with its lightweight architecture compared to ViT. Some well-established vision architectures, such as ResNet18 [24] and ViT-B [17], are overweight for online RL despite being highly acclaimed in computer vision. Online RL requires frequent data interactions to continuously update the policy, and some large-scale model architectures, even pre-trained versions, consume much wall-clock time [25, 26, 27]. Scalable ViT and the pre-trained ResNet have been validated as visual backbones for training off-policy algorithms [23, 10]. Therefore, for our work, a scaled ViT with two stems is constructed as the visual backbone to fuse RGB and depth information for online training.
II-B Generalization in Visuomotor Control
RGB has become an indispensable option when training with visual modality [1, 28]. Like human binocular perception, RGB images enable faster comprehension of complex scenes through scene changes. The sim2real discrepancy persists despite photorealistic rendering and necessary domain randomization in some simulators. A robust visual backbone with strong generalization has a crucial impact on the sim2real transfer. Specifically, an ideal visual backbone focuses on task-relevant regions. Manipulation centricity [29] is proposed as an indicator to quantify the performance of trained visual backbones for downstream manipulation tasks. Pre-training on robot-related datasets [30] can better reduce the domain gap compared to those on daily-life datasets. Similarly, related works attempt to obtain disentangled or agent-related representations that perform robustly against distracting backgrounds [11, 12]. However, some sim2real works primarily treat RGB as the main modality while rarely utilizing depth information [31, 32]. The embedded distance information and the insensitivity to appearance variations are valuable properties of depth information for enhancing visual generalization. Injection of random noise during simulation followed by a denoising process is sufficient without elaborate domain randomization for depth information [33]. In this paper, we do not simply adopt the early fusion via channel concatenation. Instead, the self-attention mechanism of ViT is used to deeply exploit the depth modality.
III Preliminaries
III-A Visual RL
The visual RL process can be formulated as a partially observable Markov decision process (POMDP), defined as a tuple $\langle\mathcal{O},\mathcal{A},\mathcal{P},\mathcal{R},\rho,\gamma\rangle$ with visual observation space $\mathcal{O}$, action space $\mathcal{A}$, transition function $\bm{o}_{t+1}=\mathcal{P}\left(\cdot|\bm{o}_{t},\bm{a}_{t}\right)$, reward function $r_{t}=\mathcal{R}\left(\bm{o}_{t},\bm{a}_{t}\right)$, distribution of the initial state $\bm{o}_{0}\sim\rho$, and discount factor $\gamma\in[0,1)$. A stack of recent frames is adopted as the current observation $\bm{o}_{t}=\left(\bm{x}_{t-i},\ldots,\bm{x}_{t}\right)$ to learn the temporal information. Our goal is training policies that maximize the cumulative reward $\mathbb{E}_{\pi}\left[\sum_{t=o}^{\infty}{\gamma}^{t}r_{t}\right]$ and finish the corresponding manipulation task.
III-B Vision Transformer
ViT is an outstanding application to extend the self-attention in natural language processing to the field of computer vision. An image $\bm{o}_{t}\in\mathbb{R}^{H\times W\times C}$ is split into $N=HW/P^{2}$ non-overlapping patches $\bm{p}_{n}\in\mathbb{R}^{N\times D}$ and expanded into a 1-dimensional sequence, where $D=P^{2}C$. The CLS token is integrated in the patch sequence, and the positional embeddings $\bm{E}_{pos}\in\mathbb{R}^{\left(N+1\right)\times D}$ are added for the preservation of the positional information
$\bm{z}_{e}^{0}=[\bm{T}_{\text{CLS}};\bm{p}_{1};\cdots;\bm{p}_{N}]+\bm{E}_{pos}.$ (1)
The transformer encoder consists of $L$ identical layers, each of which includes the multihead self-attention (MSA) and feed-forward (FFN) blocks. The input vector $\bm{z}_{e}^{i-1}\in\mathbb{R}^{N\times D},i=0\ldots L$ to MSA is first linearly mapped into query, key, and value, i.e., $\bm{Q},\bm{K},\bm{V}\in\mathbb{R}^{h\times N\times D_{h}}$. The attention score is then derived with
$\bm{O}_{ij}=\text{Softmax}\left(\frac{\bm{Q}_{ij}\bm{K}_{ij}^{T}}{\sqrt{\bm{D}_{h}}}\right)\cdot\bm{V}_{ij}\quad j=1\cdots h$ (2)
and added with $\bm{z}_{e}^{i-1}$ through the residual connection. The entire encoding process with FFN can be summarized as
$\bm{z}_{e}^{i}=$ $\ \text{MSA}\left(\text{LayerNorm}\left(\bm{z}_{e}^{i-1}\right)\right)+\bm{z}_{e}^{i-1}$
$\bm{z}_{e}^{i}=$ $\ \text{FFN}\left(\text{LayerNorm}\left(\bm{z}_{e}^{i}\right)\right)+\bm{z}_{e}^{i}$
$\bm{z}_{e}^{L}=$ $\ \text{LayerNorm}\left(\bm{z}_{e}^{L}\right).$ (3)
The CLS token of $\bm{z}_{e}^{L}$ can be treated as the visual representation $\bm{z}_{t}$ for downstream RL [23]. Masked autoencoder (MAE) [34] is a self-supervised learning scheme that builds upon ViT. Random patches are masked, and the remaining portion is utilized to construct the latent, which is then employed by the decoder to reconstruct the original image.
IV Method
This section presents a visual backbone based on ViT to fuse RGB and depth images for visual RL. The proposed visual backbone can be integrated into the popular off-policy RL framework to extract the representation that embeds multimodal visual information. To further improve the generalizability and the perception of task-related regions, a contrastive objective is formulated to maximize the mutual information between all and visible latents after random masking and encoding by ViT. Details will be given in the following sections.
IV-A Multimodal Visual Encoder
Figure 2: Overview of our approach. We build a visual backbone with two CNN stems to process RGB and depth images, respectively. The convoluted features are concatenated and then passed to ViT to obtain encoded tokens. Average pooling is applied to the unmasked tokens to derive compact visual embeddings for the policy head. Concurrently, contrastive training is performed using masked and unmasked tokens. Trained visual and policy networks are frozen and transferred to reality without fine-tuning.
ViT is commonly used to encode RGB images. For the depth stream, we refer to [35] to build another stem to split the stack of depth images, which is shown in Fig. 2. Meanwhile, prior studies have indicated that early convolutions before feeding to the ViT help to capture fine-grained details. Thus, RGB and depth images are processed separately via the corresponding convolution stems:
$\bm{h}_{t}^{rgb}=g_{conv}^{rgb}\left(\bm{o}_{t}^{rgb}\right)\qquad\bm{h}_{t}^{d}=g_{conv}^{d}\left(\bm{o}_{t}^{d}\right).$ (4)
Individual 2D sin-cos positional and 1D modality embeddings are added to the convoluted features, respectively. Embedding vectors of different modalities are concatenated and passed to ViT $\left(g_{v}\right)$. The visual representation can be obtained by the average-pooling operation on the encoded tokens without an additional MLP head:
$\bm{z}_{t}=g_{v}\left(\bm{h}_{t}\right)\quad\bm{h}_{t}=\left[\bm{h}_{t}^{rgb};\bm{h}_{t}^{d}\right].$ (5)
The encoded tokens can also be extracted after random masking and then transmitted to an asymmetric decoder to reconstruct the unmasked images of different modalities, which is the implementation of MAE. A research [35] similar to ours reconstructs RGB and tactile observations. In contrast, the effect of the asymmetric decoder on generalization will be evaluated with more challenging scenarios in Sec. V-C.
IV-B Contrastive Learning
Contrastive unsupervised learning [36] (CURL) aims to maximize the mutual information between the anchor and target. For applications in visual RL, the query and key of the latent representation can be derived by the visual encoder and span different viewpoints [37] or temporal dimensions [38]. Moreover, CURL contributes to improving the sample efficiency. Inspired by CURL and MAE, masked and unmasked image patches naturally form the anchor and target parts of CURL. Unlike MAE, random masking is performed on the convolutional features instead of normalized images and performed uniformly over two visual modalities. Random shift [22] is employed to augment the observation before random masking. Random shuffle is not executed after random masking since there is no asymmetric decoder. Note that random masking is only available to the CURL phase. The visual representation employed by the actor and critic is unmasked. The loss function of CURL can be represented as
$\mathcal{L}_{curl}=-\alpha\log\frac{\exp\left(q^{T}\cdot k^{+}/\tau\right)}{\exp\left(q^{T}\cdot k^{+}/\tau\right)+\sum_{-}\exp\left(q^{T}\cdot k^{-}/\tau\right)}$ (6)
where $\tau$ is the temperature coefficient. Query $q=\bm{z}_{t}$ and positive samples $k^{+}=g_{v}\left(\bm{h}_{t}^{mask}\right)$, where $\bm{h}_{t}^{mask}\sim p^{mask}\left(\bm{h}_{t}^{mask}|\bm{h}_{t},m\right)$. $k^{-}$ is the negative samples from different timesteps. $\alpha$ is the contrastive loss weight, and $m$ is the masking ratio. In contrast to the original CURL [36], the query-key pair is generated not by the target network for the visual encoder but instead by the recent weights.
Algorithm 1 Training details
1:Policy $\pi_{\theta}$, critics $Q_{\theta_{k}}$, visual encoder $f_{\xi}$, learning rate $\eta$, replay buffer $\mathcal{D}$, training steps $T$, evaluation period $E$, mask ratio $m$, contrastive loss weight $\alpha$, decay coefficient $\lambda$, domain randomization flag $p_{f}$, and CURL frequency $F$.
[…]
3:for timesteps $t=1\ldots T$ do
4:  $\bm{\tilde{a}}_{t}\leftarrow\pi_{\phi}\left(\cdot|\bm{o}_{t}\right)+\epsilon$
5:  $\bm{o}_{t+1}\leftarrow\mathcal{P}\left(\cdot|\bm{o}_{t},\bm{\tilde{a}}_{t}\right)$
6:  $\mathcal{D}\leftarrow\mathcal{D}\cup\left(\bm{o}_{t},\bm{\tilde{a}}_{t},r_{t},\bm{o}_{t+1}\right)$
7:  $\mathcal{B}\leftarrow\mathcal{D}$ $\qquad\vartriangleright\text{Sample batch transitions}$
[…]
10:  if $t\ \%\ F==0$ then
[…]
12:  end if
13:  if $t\ \%\ E==0$ then
14:   $p_{eval}\leftarrow$ Evaluations
15:   if $p_{eval}\geq p_{f}$ then
16:     Domain randomization enabled
17:   end if
18:  end if
19:end for
IV-C Reinforcement Learning Backbone
The proposed visual encoder and contrastive learning process are integrated into the DrQ-v2 [22] framework, which is an extension of DDPG [39] for visual RL. $N$-step returns and the clipped-double Q trick are adopted to reduce the temporal difference error and the overestimation bias of the target Q-value. Weak augmentation, i.e., random shift, is employed in DrQ-v2 to acquire the augmented data, which is limited for improving visual generalization. This paper aims to transfer the trained visual backbone and policy network to reality in a zero-shot manner. Thus, strong-augmented observations used in this paper are obtained by randomly overlaying out-of-domain images. Besides, the original data augmentation recipe in DrQ-v2 is replaced with SVEA [23] to further improve visual generalization and stabilize the training process. Given the batch data $\mathcal{B}=\left(\bm{o}_{t},\bm{a}_{t},r_{t:t+n-1},\bm{o}_{t+n}\right)$ from the replay buffer $\mathcal{D}$, the actor is updated with the following loss:
$\mathcal{L}_{\pi}=-\mathbb{E}_{\mathcal{B}\sim\mathcal{D}}\bigg{[}\min_{k=1,2}Q_{\theta_{k}}\bigl{(}f_{\xi}\left(\bm{o}_{t}\right),\bm{\tilde{a}}_{t}\bigr{)}\bigg{]},$ (7)
where $f_{\xi}$ indicates the vision encoder. $\bm{\tilde{a}}_{t}=\pi_{\phi}\left(f_{\xi}\left(\bm{o}_{t}\right)\right)+\epsilon$ is inferred by the policy network only with the weak augmented stream. $\epsilon$ is the exploration noise. $\bm{o}_{t}^{\text{aug}}$ denotes the observation after the strong augmentation. The critic loss is defined with both weak and strong augmented streams:
$\mathcal{L}_{Q}=$ $\mathbb{E}_{\mathcal{B}\sim\mathcal{D}}\biggl{[}\beta_{1}\left\lVert Q_{\theta_{k}}\left(f_{\xi}\left(\bm{o}_{t}\right),\bm{a}_{t}\right)-y_{t}^{tgt}\right\rVert_{2}^{2}+$ (8)
$\beta_{2}$ $\left\lVert Q_{\theta_{k}}\left(f_{\xi}\left(\bm{o}_{t}^{\text{aug}}\right),\bm{a}_{t}\right)-y_{t}^{tgt}\right\rVert_{2}^{2}\biggr{]}\quad\forall k\in\left\{1,2\right\},$
where
$y^{tgt}=\sum_{i=0}^{n-1}\gamma^{i}r_{t+i}+\gamma^{n}\bigg{(}\underset{k=1,2}{\rm{min}}Q_{\bar{\theta}_{k}}\left(f_{\xi}\left(\bm{o}_{t+n}\right),\bm{\tilde{a}}_{t+n}\right)\bigg{)}$ (9)
is the critic target obtained only with unaugmented observation. $\beta_{1}$ and $\beta_{2}$ are constants utilized to balance the two data streams.
Figure 3: Benchmark environments and unseen scenarios for agents during testing. (Top) Three manipulation tasks are labeled from left to right as Assembly, Lift, and PickAndPlace. (Bottom) Challenging color hard environments.
IV-D Curriculum-based Domain Randomization
Domain randomization is incorporated into the training process to reduce the sim2real gap, typically encompassing variations in scene appearance, lighting conditions, dynamic parameters, and camera viewpoints. Large randomized magnitude will trigger divergence in the RL process. Thus, we implement a curriculum-based domain randomization approach inspired by [33] to progressively broaden the random range. When domain randomization initiates, the random ranges of all parameters expand exponentially with increasing episodes:
$R_{c}=R_{d}\cdot\lambda+R_{r}\cdot\left(1-\lambda^{T_{e}}\right),$ (10)
where $R_{c}$ and $R_{d}$ are randomized and default domain parameters, respectively. $R_{r}$ represents the predefined random range. $\lambda$ is the decay coefficient over running episodes $T_{e}$. However, domain randomization is invoked when a fixed number of training frames is achieved in [33], which is an empirical determinant for different tasks. A more flexible determinant is adopted in this paper, i.e., the evaluation success rate. Such a metric prevents training divergence and unnecessary computational overhead caused by either premature or delayed initiation of domain randomization. Algorithm 1 provides a detailed explanation about the training process.
Figure 4: Training curves using different fusion approaches, where solid lines and shaded areas represent the mean and confidence interval over 5 random seeds, respectively.
V Simulations
This section begins with presenting manipulation tasks and simulation settings used for subsequent testing. Next, different fusion methods for RGB and depth are tested and quantitatively analyzed on the above tasks to serve as references for subsequent sim2real transfer. Last, ablation experiments are finally performed to verify the effect of some potential factors.
V-A Tasks
We develop three manipulation tasks powered by MuJoCo as benchmarks for evaluations, which are shown in Fig. 3. For each task, the KUKA LBR IIWA 14 R820 manipulator is adopted as the main embodiment, and stage-dense rewards are customized. Object position is randomized during episode resets. The details of all tasks are as follows:
Assembly: Modular connectors mounted on the end-effector and table are genderless and serve for on-orbit assembly. The collision meshes of connectors are manually segmented before import to optimally maintain their surface contours. This task is considered successful when both connectors achieve the mating state within a tolerance of $1cm$. The agent is endowed with 3-DOF motion in the Cartesian space. Due to the presence of numerous contact-rich states during assembly, we extend the action space with variable impedance gains [40].
Lift: The Robotiq 2F-85 gripper is attached to the end-effector to grasp the red cube. The action space also contains 3-DOF motion in the Cartesian space with an additional continuous control of the gripper. The red cube should be lifted above the table.
PickAndPlace: This task has the same action space as Lift. Episodes are determined to be successful by picking the red cube first and then placing it in the yellow box.
TABLE I: Generalization in unseen environments. Success rate (mean $\pm$ standard deviation) is evaluated over $100$ random episodes.
Method Assy. (%) Lift (%) P&P (%)
Ours $\bm{82.8\pm 1.8}$ $\bm{39.8\pm 1.7}$ $\bm{42.0\pm 4.4}$
EF $61.8\pm 13.1$ $14.8\pm 6.1$ $5.4\pm 4.5$
LF $77.4\pm 8.6$ $26.0\pm 10.5$ $14\pm 7.4$
TABLE II: Grad-CAM visualization of different visual backbones on all tasks.
Task Ours EF LF
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
V-B Simulation Results
This section compares our approach with different fusion approaches: (i) Early fusion (EF) is realized using channel-wise concatenation in a CNN-based encoder, followed by ReLU activation [22]. (ii) Late fusion (LF) doubles the CNN stem in the EF scenario to process different modalities separately. Different encoded representations are flattened and then concatenated before being passed to the policy head.
A maximum depth value is set to normalize the depth image and remap to the pixel range [0, 255]. Additional Gaussian and depth-independent noises are added and smoothed with Gaussian blur [33]. The visual observation consists of $3$ RGB frames and $1$ depth frame, each with a size of $128\times 128$. Our visual backbone includes a 4-layer ViT with 8 heads and 2 convolutional stems with a 4-layer CNN. All hyperparameters for training are detailed in Appendix.
Figure 5: Attention visualization of RGB (left), depth (middle), and RGB-D (right) with the proposed visual backbone.
Five runs are performed with different seeds for each method, and the training processes without domain randomization are shown in Fig. 4. Our method demonstrates better sample efficiency, particularly in the PickAndPlace task. Generalization tests are conducted in the unseen scenarios illustrated in Fig. 3. The test results in Tab. I are obtained through 5 distinct model weights, each trained on a limited number of frames illustrated in Fig. 4. Meanwhile, Grad-CAM [41] visualizations are presented in Tab. II to analyze the underlying effects of visual backbones on generalization performance. Compared to EF and LF, our visual backbone shows broader attention on agent- and object-related regions. More task-centric areas can be captured and extracted as informative representations for downstream inference, which contributes to the improved performance of our method in Tab. I. Meanwhile, the complementary properties of two modalities are exploited by our visual backbone to focus on different task-related regions, as illustrated in Fig. 5. For EF and LF, the dual-CNN backbone outperforms channel concatenation for multimodal feature fusion.
Figure 6: Image reconstruction with the asymmetric decoder.
TABLE III: Ablation experiments with the proposed visual backbone on the lift task.
Ablation Success rate (%)
w/ CURL $39.8\pm 1.7$
w/o. CURL $24.6\pm 4.5$
w/ decoder & w/o. CURL $31.8\pm 6.11$
Figure 7: Real-world setup for tasks Assembly, Lift, and PickAndPlace.
TABLE IV: Sim2Real results with the proposed visual backbone.
Tasks Standard (%) Challenging (%)
Assy. $15/15$ $15/15$
Lift $14/15$ $13/15$
P&P $14/15$ $12/15$
V-C Ablations
The following ablation experiments are performed on the Lift task under the condition of 400K training frames.
CURL. The impact of CURL on generalization is explored via ablation experiments. Eq. 6 is disabled, and the visual backbone is only updated via Eq. 8. The success rate on the Lift task decreased by approximately $15\%$, demonstrating that the proposed CURL scheme effectively enhances the generalization of the encoder.
Figure 8: Zero-shot transfer with the proposed visual backbone in standard (left) and challenging (right) real-world scenarios.
Decoder. Masked tokens are passed to a 3-layer, 4-head decoder to reconstruct the original images [34, 35]. Ablation studies are performed to investigate the effect of decoders on generalization without CURL. The decoder is updated with the following loss:
$\mathcal{L}_{dec}=w_{rgb}\cdot\text{MSE}\left(\bm{p}_{rgb},\hat{\bm{p}}_{rgb}\right)+w_{d}\cdot\text{MSE}\left(\bm{p}_{d},\hat{\bm{p}}_{d}\right),$ (11)
where $\bm{p}_{rgb}$ and $\bm{p}_{d}$ represent RGB and depth image patches without convolutional processing in Fig. 2. $\hat{\bm{p}}_{rgb}$ and $\hat{\bm{p}}_{d}$ are patches predicted by the decoder. $w_{rgb}=1$ and $w_{d}=10$ are reconstruction weights for different modalities. $\mathcal{L}_{dec}$ is added to Eq. 8 and updated along with the critic. Fig. 6 shows the predicted RGB and depth images at a masking ratio of $50\%$. With the current hyperparameter settings, the decoder can ignore task-irrelevant backgrounds and exhibit potential to improve the generalization compared to the case without CURL, as presented in Tab. III.
VI Sim2real Experiments
The KUKA robot controller, PC server, and PC client construct the real-world communication network via Socket. The PC server bridges the other two parts and delivers visual information from the RealSense L515 to the PC client. Additional communication is required to achieve the gripper control via the PC server. The trained policy is deployed on the PC client and makes inferences based on the received data. Fig. 7 depicts the experimental setup, which inherits the basic deployment of simulation in Fig. 3. Domain randomization covers scene appearance, camera position, and lighting position. End-effector control is activated for all tasks, and dynamics randomization is excluded [40]. The training frames are expanded to 800K before the sim2real transfer.
The trained policy is directly transferred to the experimental platform and tested with 30 episodes, half of which are constructed in visually challenging environments. We increase the difficulty of visual observation by adding visual distractors, unseen backgrounds, and dark light on the basis of standard scenarios. Tab. IV summarizes experimental results and demonstrates that our method is effective for transfer in a zero-shot manner. Meanwhile, Fig. 8 illustrates execution processes in different visual environments using our method. The trained policy enables the KUKA arm to complete the assembly of genderless connectors with variable impedance gains to achieve desired fitting clearance. The material of the genderless connectors makes it difficult to replicate the reflective visual effect in simulation. The experiment results also confirm that our visual backbone and policy remain robust to discrepancies in material or scene appearance after the training process incorporating data augmentation and curriculum-based domain randomization. Moreover, for visually challenging scenarios, RGB exhibits large discrepancies compared to those in simulation. Here, our visual backbone can unleash the power of depth information and extract generalized visual representations to complete the task, which is essential for object-grasping tasks.
VII Conclusion And Limitations
This paper focused on exploring the effect of depth modality and visual backbones on generalization and sim2real performance of visual reinforcement learning policies. A visual backbone integrating ViT and CNN is first presented to fuse RGB and depth modalities. A contrastive unsupervised learning scheme is then constructed to extract consistent information from masked and unmasked tokens processed by ViT. The simulation results show that the fusion of depth information has a direct impact on the generalization of visual reinforcement learning policies via different visual backbones, and our proposed backbone demonstrates better performance. A curriculum-based domain randomization is developed for sim2real transfer. Real-world experiments indicate that our policy can realize the sim2real transfer in a zero-shot manner and adapt to diverse visual scenarios.
The early and late fusion baselines in this paper are constructed based on common CNN architectures. However, numerous ConvNets remain available for comparison. This work demonstrates that ViT-based backbones are promising alternatives to some ConvNets in visual RL and sim2real applications. We are also dedicated to exploring the potential of transformer-like architectures, which is consistent with many prior studies in computer vision [42, 43]. Besides, further work will dive into the generalization to different camera viewpoints.
Appendix A Hyperparameters
TABLE V: Hyperparameters used for training
Hyperparameters Value
Image size 128 $\times$ 128
Frame stack 3 (RGB), 1 (Depth)
Discount factor $\gamma$ 0.99
Replay Buffer size 1e6
Batch size 256
N-step return 3
Optimizer Adam
Learning Rate 1e-4
Max episode steps 500 (PickAndPlace), 400 (Others)
Exploration schedule Linear $\left(1.0,0.1,100\text{K}\right)$ (Others)
Linear $\left(1.0,0.1,300\text{K}\right)$
Action repeat 2
CNN encoder Conv (c=[32, 64, 128, 256])
Actor & Critic Linear (c=[256, 1024, 1024])
Patch size of ViT 16
Heads of ViT 8
Embedding dim of ViT 512
Head dim of ViT 32
Layers of ViT 4
Mask ratio for CURL 0.1